OpenStudy (anonymous):

somebody can help me to integrate this? 2^x /(4+4^x) dx

OpenStudy (tkhunny):

Have you noticed that \(4^{x} = 2^{2x} = \left(2^{x}\right)^{2}\)?

OpenStudy (dumbcow):

lets do some substitutions \[u = 2^x\] \[du = \ln(2) 2^x dx\] the 2^x on top will cancel leaving \[\frac{1}{\ln 2} \int\limits \frac{du}{4 + u^2}\] now use identity --> 1 +tan^2 = sec^2, to help make next substitution \[u = 2 \tan \theta\] \[du = 2 \sec^2 \theta\] \[\rightarrow \frac{1}{\ln 2} \int\limits \frac{2 \sec^2 \theta}{4 + 4\tan^2 \theta}\] from identity, the sec^2 will cancel leaving \[\frac{1}{2 \ln 2} \int\limits d \theta\] \[= \frac{1}{2 \ln 2} \theta = \frac{1}{2 \ln 2} \tan^{-1} \frac{u}{2}\] \[=\frac{1}{2 \ln 2} \tan^{-1} (\frac{2^x}{2})\]